Suppose a linear production with LaTeX: Y\:=\:100N Y = 100 N and a government that consumes 5 units of consumption goods. Suppose that the economy is in equilibrium and the optimal choice of consumption is equal to 20, then the optimal choice of leisure is

a. 0.75

b. 0.1

c. 0.5

d. 0.9
Answer #1

we have Y = C + G

Now, in equilibrium, C = 20. G is always = 5

Thus, Y = 20 + 5 = 25

We have, Y = 100N

N = Y/100

Thus N = 25/100 = 0.25

Hence, optimal choice of leisure = 1 - choice of labor = 1-0.25 = 0.75